# Refinable modeling for unbinned SMEFT analyses

### v1

@@ Abstract
We present techniques for estimating the effects of systematic uncertainties in unbinned data analyses at the LHC. Our primary focus is constraining the Wilson coefficients in the standard model effective field theory (SMEFT), but the methodology applies to broader parametric models of phenomena beyond the standard model (BSM). We elevate the well-established procedures for binned Poisson counting experiments to the unbinned case by utilizing machine-learned surrogates of the likelihood ratio. This approach can be applied to various theoretical, modeling, and experimental uncertainties. By establishing a common statistical framework for BSM and systematic effects, we lay the groundwork for future unbinned analyses at the LHC. Additionally, we introduce a novel tree-boosting algorithm capable of learning highly accurate parameterizations of systematic effects. This algorithm extends the existing toolkit with a versatile and robust alternative. We demonstrate our approach using the example of an SMEFT interpretation of highly energetic top quark pair production in proton-proton collisions.
@@ 1 Introduction
The Large Hadron Collider (LHC) is a veritable gold mine of data, where complex detector signatures emanate from the decays of particles produced in hard-scatter interactions at the highest achievable energy scales. It presents a challenge and a fertile ground for exploring fundamental physics. Recent developments in machine learning (ML) algorithms provide a rich technological toolkit that goes far beyond improving reconstruction and object-tagging performances but opens the door to novel analysis techniques. High-dimensional unbinned data analyses, where the predictions for dozens of features probe an equally large number of model parameters, become feasible via machine-learned surrogates trained to provide optimality in the context of the underlying hypothesis tests.
In the theoretical domain, the absence of any compelling signal of new resonant phenomena fuelled the rise of the standard model effective field theory (SMEFT) [1, 2, 3, 4, 5] as the preferred language for describing hypothetical phenomena below an assumed energy scale chosen at $\Lambda_{\textrm{SMEFT}}=1$  TeV and separating the LHC energy scale from resonant phenomena at the BSM energy scale. The SMEFT is a powerful framework extending the standard model (SM) Lagrangian by field monomials whose prefactors, the Wilson coefficients, are the parameters of interest (POIs) in experimental measurements. These give the experimentalists access to a wide range of hypothetical high-scale models without dealing with the complexity of those models’ fundamental parameters. The SMEFT organizing principle is the operators’ mass dimension, starting at six for the new physics scenarios relevant at the LHC [6]. Because the Wilson coefficients’ lowest-order matrix-element (ME) modifications are linear, we can accurately describe the cross-section deviations by quadratic polynomials within the SMEFT’s range of validity [7]. This simple analytic structure is exploited in various simulation-based inference techniques that forecast significant performance gains, in particular when many Wilson coefficients are probed simultaneously [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19]. Such algorithms provide statistically optimal observables at the detector level, promising optimal performance and fast evaluation after an upfront training stage.
@@
Nevertheless, most current LHC measurements are straightforward Poisson counting experiments, partially because these reduce the computational demand. Large-scale computing infrastructure has become more accessible, but the application of unbinned techniques is still hampered by the absence of a comprehensive set of tools that bring decades of experience with treating systematic effects in binned Poisson measurements on par with the unbinned case. The available methodology for treating systematic uncertainties in unbinned simulation-based inference techniques, such as ML optimal observables, is sparse. While optimal ML observables have a sound footing in well-developed statistical methodology, the otherwise finely honed procedures for treating systematic uncertainties are rarely seen in this light. The present work aims to change that situation through a comprehensive statistical interpretation of procedures for treating systematic effects. We explain how the factorization of individual systematic effects facilitates the training of multi-variate parametrized regressors and how to address uncertainties in the normalization of processes. The most significant advantage of this approach is its stage-wise nature. Adding new processes or systematic uncertainties does not invalidate partially available training.
@@
Figure 1: A hierarchy of scales governs our assumed understanding of physical phenomena and the modeling of collider physics. Staged event simulation proceeds in the “forward mode” from right to left, while data analyses proceed in the inverse direction, aimed at constraining models of UV phenomena or the SMEFT POIs from the detector-level measurements. The sketch is adapted from Ref. [20].
@@
The conceptual cornerstone of the modeling of collider phenomena, depicted in Fig. 1, is the hierarchical separation of phenomena according to their energy scale, starting with the landscape of hypothetical UV phenomena and their SMEFT parametrization at the scale of $\Lambda_{\textrm{SMEFT}}$. We aim at parameter inference through frequentist confidence intervals [21] in the space of the SMEFT Wilson coefficients, our POIs, denoted by ${\boldsymbol{\theta}}$. Those parameters are, therefore, not stochastic. However, we note that there is no conceptual limitation to Bayesian SMEFT analysis.
We reflect the hierarchy in Fig. 1 by grouping unobserved (latent) variables and the accompanying systematic effects into the parton, particle, and detector levels. This somewhat arbitrary choice is a reasonable compromise of the need for sufficient detail and an acceptable level of notational clutter; more levels can be introduced, but the ensuing complexity would disguise the simple concepts at work. This staging is also paralleled in collider event simulation. At the parton level, ME generators sample the SMEFT ME-squared terms. Important systematic uncertainties at this stage pertain to the unphysical effects of technically unavoidable energy scales related to the perturbative renormalization of the fixed-order predictions and the factorization of collinear and infrared radiation. Uncertainties in the parton distribution functions (PDFs) cover our lack of knowledge in the composition of our initial state, that is, $\mathrm{p}\mathrm{p}$-collisions at the LHC. Proceeding from $\Lambda_{\textrm{SMEFT}}$ to $\Lambda_{\textrm{QCD}}$, we reach the particle level: dedicated computer codes simulate effects from the parton shower (PS), the fragmentation of strongly interacting particles, their hadronization, the decay of the hadrons, and the modeling of the underlying event. The tuning of partially empirical models is subject to a plethora of systematic uncertainties related to those models’ parameters. Beyond tuning, treating the collective differences among competing models or alternative generators as a “two-point” systematic uncertainty can be desirable. Finally, the detector level comprises the simulation of the particles’ interaction with the detector material, the digitization, the trigger logic, and the event reconstruction, including shortcomings and bugs in the computer code. We attempt to provide the generic recipes for obtaining ML parametrizations of all these effects by grouping them into equivalent classes of systematic variations that we tackle individually. With the procedures in place, we explain how to build unbinned models in a refinable manner. Accounting for new effects or background contributions can proceed iteratively, and the surrogates trained for an existing model are not invalidated when upon its extension.
@@
On the technical side, we also close a gap in the methodological toolbox by developing a tree-boosting algorithm capable of learning arbitrarily accurate parameterizations of systematic effects. This is achieved by extending the predictive capacity of tree algorithms to provide regressors that are parametric in externally provided data; in our case, these are the nuisance parameters associated with the systematic effects. The ensuing “Parametric regression tree” is a robust and versatile alternative to neural networks for this application. Its training setting entirely rests on the statistical foundation of unbinned model building. As a side-effect, this allows one to understand tree-boosting entirely in the context of unbinned hypothesis tests. The trees’ terminal nodes can be seen as measurement bins where we can inject an analytic dependence on the nuisance parameters equivalent to the standard binned case. The parametric regression tree thus learns an arbitrary expressive surrogate of differential cross-section ratios (DCRs) in terms of a possibly high-dimensional set of model parameters.
Throughout the developments, unbinned SMEFT analyses serve as the motivating use case, and we assume throughout that the POI dependence is learned by one of the algorithms available in the literature [8, 9, 10, 11, 12, 13, 15, 16, 17, 14]. We demonstrate the application of our tools for learning systematic effects in a simple semi-realistic case study of SMEFT constraints in top quark pair production.
@@
The rest of the paper is structured as follows. We discuss the relation to existing works for unbinned (SMEFT) analyses in Sec. 2. The statistical setup for unbinned hypothesis tests is provided in Sec. 3. A comprehensive and elementary review of the statistical interpretation of event simulation, suitable for developing ML tools, is given in Sec. 4. This part also defines the terminology for Sec. 5, which explains how to train generic regressors for suitable parametrizations of the various model-parameter dependencies. In Sec. 6, we use such tools as building blocks for constructing refinable unbinned models. The parametric regression tree is built in Sec. 7, and a comprehensive interpretation of the various algorithmic steps is provided. The new tools are applied in Sec. 8, where we demonstrate the application of the procedures for SMEFT analyses of top quark pair production in the two-lepton channel. We provide conclusions in Sec. 9.
@@ 2 Relation to other works
The Boosted Information Tree (BIT[10, 9] is a tree-based algorithm for learning SMEFT effects. In this work, we adopt it for the SMEFT signal modeling and extend it toward predicting the full positive quadratic SMEFT polynomial. The BIT algorithm learns the quadratic SMEFT polynomial using the same statistical foundation as the present work. The parametric regression tree in Sec. 7 has a different goal, but the technical implementation and the statistical interpretation of the boosting are closely related.
@@
The authors of the Madminer framework [13, 14, 15, 16, 17, 18] developed the understanding of event simulation for likelihood-free inference that is also an essential basis for this work. On the technological side, Madminer provides various techniques for general parameter inference and, specifically, also for unbinned SMEFT analyses. Beyond these motivating examples, Madminer arguably established the simulation-based inference methods as a subfield in high-energy physics, to which the present work contributes. While Madminer can also model systematic uncertainties, we use a more general and incrementally refinable statistical model for this purpose.
@@ 3 Unbinned likelihood ratio tests
When a data set $\mathcal{D}$ is available, confidence level (CL) intervals associated with the POIs ${{\boldsymbol{\theta}}}$ and nuisance parameters ${\boldsymbol{\nu}}$ can be determined from the profiled likelihood ratio test statistic
@@
Here, $\mathcal{D}$ is our primary data, while $\mathcal{A}$ denotes an auxiliary data set, possibly comprising several non-overlapping components, each designed for constraining different systematic uncertainties such as the luminosity or the calibration of reconstructed jet momenta. We denote the maximum-likelihood estimate (MLE) of the nuisance parameters for a given ${{\boldsymbol{\theta}}}$ by $\boldsymbol{\hat{\nu}}_{{\boldsymbol{\theta}}}$. The pair $(\boldsymbol{\hat{\theta}},\boldsymbol{\hat{\nu}})$ is the MLE obtained simultaneously for all the model parameters. By construction, $q_{{{\boldsymbol{\theta}}}}(\mathcal{D})$ is semi-positive definite for any value of ${{\boldsymbol{\theta}}}$, and larger values indicate increasing incompatibility between the model family described by ${{\boldsymbol{\theta}}}$ and the data set $\mathcal{D}$. In the absence of nuisance parameters, the Neyman-Pearson lemma states that a hypothesis test of fixed size $\alpha$ based on $q_{\boldsymbol{\theta}}(\mathcal{D})$ has the highest power, that is, it has the maximal probability to correctly reject the null-hypothesis when the alternate is true [25].
Technically, $\mathcal{A}$ should appear as an argument of $q_{{{\boldsymbol{\theta}}}}(\mathcal{D})$, but we drop it in the notation because an analytic approximation of the corresponding factor in the likelihood, described below, will encode all its effects. We assume that $\mathcal{A}$ and $\mathcal{D}$ do not overlap and that the SMEFT effects, parametrized by ${\boldsymbol{\theta}}$, are negligible in $\mathcal{A}$. The auxiliary data set then produces a multiplicative term $L(\mathcal{A}|{\boldsymbol{\nu}})$ in the likelihood,
@@
According to a powerful result by Wilks [26], if $\mathcal{D}$ are distributed under $\boldsymbol{\theta}$, then $q_{{{\boldsymbol{\theta}}}}(\mathcal{D})$ is asymptotically distributed as a $\chi^{2}_{N_{{{\boldsymbol{\theta}}}}}$ distribution, where $N_{{{\boldsymbol{\theta}}}}$ is the number of independent degrees of freedom in ${{\boldsymbol{\theta}}}$. This asymptotic distribution is, therefore, independent of the value of the nuisance parameters. The ensuing simplification in the limit-setting procedure is why the LHC data analyses almost exclusively adopt the profiled likelihood ratio test statistic. The confidence intervals at, e.g., a level of $1-\alpha=95\%$ are then given by the solution of
@@
where $F_{\chi^{2}_{N_{\boldsymbol{\theta}}}}$ is the cumulative distribution function of the $\chi^{2}_{N_{{{\boldsymbol{\theta}}}}}$ distribution. To obtain confidence intervals with some of the POIs profiled, those parameters are treated as nuisance parameters in $q_{\boldsymbol{\theta}}$ and $N_{{\boldsymbol{\theta}}}$ reduces accordingly.
The likelihood functions in Eq. 1 are “extended” likelihoods: a counting variable with Poisson distribution describes the random fluctuation in the total number of observed events. The remainder of the discriminating information is encoded in the fiducial (detector level) probability density function (pdf) $p(\boldsymbol{x}|{{\boldsymbol{\theta}}},{\boldsymbol{\nu}})$, which relates to the fiducial differential cross section as
@@
We denote the inclusive fiducial cross-section by $\sigma({{\boldsymbol{\theta}}},{\boldsymbol{\nu}})$. In general, ${\textrm{d}}\Sigma(\boldsymbol{x}|{{\boldsymbol{\theta}}},{\boldsymbol{\nu}})$ receives contribution from several processes. With the luminosity $\mathcal{L}({\boldsymbol{\nu}})$, subject to systematic uncertainties whose effects we parametrize by nuisance parameters, the likelihood to observe a data set $\mathcal{D}$ of size $N$ can, in general, be written in terms of the differential cross-section as
@@
where the Poisson distribution with mean $\lambda$ is denoted by $\textrm{P}_{\lambda}$. The extended log-likelihood ratio corresponding to two sets of model parameters becomes
@@
Each source of the $K$ systematic uncertainties is associated with a nuisance parameter $\nu_{k}$, collectively denoted by ${\boldsymbol{\nu}}$. A plethora of systematic effects related to the detector calibration, the luminosity measurement, the theoretical computations, etc., are modeled with nuisance parameters. Measurements can constrain some of these uncertainties in the observed auxiliary data set $\mathcal{A}_{0}$, i.e., the particular instance among all possible $\mathcal{A}$ obtained in measurement. We chose ${\boldsymbol{\nu}}={\mathbf{0}}$ to correspond to the maximum of $L(\mathcal{A}_{0}|{\boldsymbol{\nu}})$ so that by definition
@@
The central value ${\boldsymbol{\nu}}={\mathbf{0}}$ thus amounts to the best available calibrations in all modeling aspects before $\mathcal{D}$ is considered. The nuisances, in turn, parametrize deviations from this hypothesis; it is not a random quantity because $\mathcal{A}_{0}$ corresponds to a single outcome of the auxiliary measurement, the one observed in data [22]. Together with ${{\boldsymbol{\theta}}}={\mathbf{0}}$, the choice defines a SM reference hypothesis with likelihood
@@
which describes the SM without any SMEFT effects and includes the best available calibrations. We can furthermore parametrize the systematic effects to make the $\nu_{k}$ uncorrelated among each other as far as possible and scale them such that the auxiliary log-likelihood ratio becomes a simple analytic expression in terms of ${\boldsymbol{\nu}}$. In the Gaussian approximation, this can be achieved by diagonalizing the Hessian of the auxiliary likelihood function at ${\boldsymbol{\nu}}={\mathbf{0}}$. In this case, we have a penalty of the form
@@
although generalizations to other pdfs are straightforward. Suppose a specific nuisance parameter should only be constrained by the primary data set but not by $\mathcal{A}_{0}$. In that case, it is excluded from the penalty term in Eq. 9 and called “floating”. While many uncertainties, e.g., related to PDFs, have a well-defined interpretation in terms of the parameters of the SM, this is not always the case. For example, uncertainties related to the renormalization or factorization scales aim to account for an unavoidable deficit of perturbative accuracy, but there is no guarantee of statistical coverage when varying those scales in the simulation. With this caveat in mind, we nevertheless treat all systematic uncertainties heuristically in the same way.
Finally, we normalize the likelihoods in Eq. 1 by separately dividing the numerator and denominator by the SM reference likelihood in Eq. 8 and find
$q_{{{\boldsymbol{\theta}}}}(\mathcal{D})=\max_{{\boldsymbol{\nu}}}u(\mathcal{D},\mathcal{A}|{\boldsymbol{\nu}},{{\boldsymbol{\theta}}})-\max_{{\boldsymbol{\nu}},{{\boldsymbol{\theta}}}}u(\mathcal{D},\mathcal{A}|{\boldsymbol{\nu}},{{\boldsymbol{\theta}}})$ (10)
@@
The main drawback of the unbinned likelihood ratio test statistic in Eq. 11 is the necessity to evaluate the DCR, the inclusive cross-section, and the luminosity parametrically as a function of the model parameters. While log-normal (multiplicative) nuisances accurately model the parameter dependence of the luminosity around the central value [27, 28] measured in $\mathcal{A}_{0}$, we also need an estimate of the DCR and the inclusive cross-section.
This cannot be achieved parametrically in the model parameters using event generators alone, as they operate in “forward” mode, i.e., as a sequence of staged stochastic processes. Event generators typically cannot estimate the DCR as a function of an externally provided simulated or real event. However, the minimization in Eq. 1 must be feasible for inference, necessitating parametric estimates from other sources. The simulation-based inference approach aims to provide the required estimates using machine learning surrogates trained on a limited number of simulated samples. In the following chapters, we will present techniques to minimize CPU-intensive simulation requirements and demonstrate how simulation-based inference methods can benefit from efficiently generated “synthetic” data sets, thereby facilitating the primary inference problem.

### v3

@@ Abstract
We present methods to estimate systematic uncertainties in unbinned LHC data analyses, focusing on constraining Wilson coefficients in the standard model effective field theory (SMEFT). Our approach also applies to broader parametric models of non-resonant phenomena beyond the standard model (BSM). By using machine-learned surrogates of the likelihood ratio, we extend well-established procedures from binned Poisson counting experiments to the unbinned case. This framework handles various theoretical, modeling, and experimental uncertainties, laying the foundation for future unbinned analyses at the LHC. We also introduce a tree-boosting algorithm that learns precise parametrizations of systematic effects, providing a robust, flexible alternative to neural networks for modeling systematics. We demonstrate this approach with an SMEFT analysis of highly energetic top quark pair production in proton-proton collisions.
@@ 1 Introduction
The Large Hadron Collider (LHC) generates vast amounts of data from particle decays in high-energy interactions, offering a unique opportunity to explore fundamental physics. Recent advances in machine learning (ML) provide powerful tools not only for reconstruction and object-tagging but also for novel analysis techniques. High-dimensional unbinned analyses, where dozens of features probe a large number of model parameters, are now feasible with machine-learned surrogates optimized for hypothesis testing.
In the theoretical domain, the lack of new resonance signals has led to the adoption of the standard model effective field theory (SMEFT) [1, 2, 3, 4, 5] as the main framework for describing phenomena below an assumed energy scale, conventionally set at $\Lambda_{\textrm{SMEFT}}=1$  TeV. This framework extends the standard model (SM) Lagrangian with field monomials, where the Wilson coefficients serve as the parameters of interest (POIs). The SMEFT enables experimentalists to test a range of high-scale models without dealing with their fundamental parameters.
The SMEFT is organized by the mass dimension of operators, beginning with dimension six for relevant new physics scenarios at the LHC [6]. Since the lowest-order matrix-element (ME) modifications to Wilson coefficients are linear, cross-section deviations can be described by quadratic polynomials within the SMEFT’s validity range [7]. This analytic structure supports simulation-based inference (SBI) methods that improve performance, especially when probing multiple Wilson coefficients simultaneously [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19]. These methods offer statistically optimal observables at the detector level, with fast evaluation after an initial training stage.
@@
Nevertheless, most current LHC measurements are straightforward Poisson counting experiments, partially because these reduce the computational demand. Large-scale computing infrastructure has become more accessible, but the application of unbinned techniques is still hampered by the absence of a comprehensive set of tools that bring decades of experience with treating systematic effects in binned Poisson measurements on par with the unbinned case. The available methodology for treating systematic uncertainties in unbinned simulation-based inference techniques, such as ML optimal observables, is sparse. While optimal ML observables have a sound footing in well-developed statistical methodology, the otherwise finely honed procedures for treating systematic uncertainties are rarely seen in this light. The present work aims to change that situation through a comprehensive statistical interpretation of procedures for treating systematic effects in SBI. We explain how the factorization of individual systematic effects facilitates the training of multi-variate parametrized regressors and how to address uncertainties in the normalization of processes. The most significant advantage of this approach is its stage-wise nature. Adding new processes or systematic uncertainties does not invalidate partially available training.
@@
Figure 1: A hierarchy of scales governs our assumed understanding of physical phenomena and the modeling of collider physics. Staged event simulation proceeds in the “forward mode” from right to left, while data analyses proceed in the reverse direction, aimed at constraining models of UV phenomena or the SMEFT POIs from the detector-level measurements. The sketch is adapted from Ref. [20].
@@
The conceptual cornerstone of the modeling of collider phenomena underpinning SBI relies on a hierarchical separation of processes by energy scale, starting from hypothetical UV phenomena and their SMEFT parametrization at $\Lambda_{\textrm{SMEFT}}$. The SMEFT Wilson coefficients, our POIs, are denoted by ${\boldsymbol{\theta}}$, and we aim at parameter inference through frequentist confidence intervals [21]. Those parameters are, therefore, not stochastic. However, we note that there is no conceptual limitation to Bayesian SMEFT analysis.
Figure 1 reflects this hierarchy, from right to left, by grouping unobserved (latent) variables and systematic effects at the parton, particle, and detector levels. This division balances sufficient detail with manageable notation complexity; more levels could be added but would obscure the core concepts. Collider event simulation mirrors this staging: at the parton level, ME generators sample SMEFT ME-squared terms. Key systematic uncertainties here include unphysical effects from technically unavoidable energy scales linked to the perturbative renormalization of fixed-order predictions and the factorization of collinear and infrared radiation. Uncertainties in the parton distribution functions (PDFs) cover our lack of knowledge in the composition of the pp initial state. At the scale of $\Lambda_{\textrm{QCD}}$, particle-level simulations handle parton shower (PS) effects, fragmentation, hadronization, decay, and underlying event modeling. Tuning these models introduces systematic uncertainties, and modeling generator differences as a “two-point” systematic uncertainty can be effective. At the detector level, simulations include particle interactions with detector material, digitization, trigger logic, and event reconstruction.
Our approach provides a general method to obtain ML-based parametrizations of all these effects by grouping them into classes of systematic variations addressed individually. With these procedures in place, we can build unbinned models that can be iteratively refined. Adding new effects or background contributions does not invalidate surrogates trained on the initial model.
@@
On the technical side, we fill a gap in the methodology by developing a tree-boosting algorithm that can learn arbitrarily accurate parametrizations of systematic effects. This is done by extending tree algorithms to produce regressors that are parametric in externally provided data; in our case, the nuisance parameters (${\boldsymbol{\nu}}$) linked to systematic effects. The resulting “Boosted Parametric Tree” (BPT) offers a robust and flexible alternative to neural networks for this purpose, with the training procedure fully grounded in unbinned model building. Therefore, it enables a full understanding of tree-boosting within the context of unbinned hypothesis tests. The terminal nodes of the trees act as measurement bins, allowing for an analytic dependence on nuisance parameters similar to the binned case. The BPT thus learns an expressive surrogate for differential cross-section ratios (DCRs), accommodating a potentially high-dimensional set of model parameters.
We use unbinned SMEFT analyses as our motivating case, assuming that the dependence on the POIs is learned by an algorithm from the literature [8, 9, 10, 11, 12, 13, 15, 16, 17, 14]. We demonstrate our tools for modeling systematic effects through a semi-realistic SMEFT case study in top quark pair production. However, this methodology is broadly applicable and could enhance the inference of any SM parameter with a non-resonant impact on collider data. In addition to extracting parameters like $\alpha_{s}$, electro-weak precision observables, or $\sin^{2}\mathbf{\theta}_{\textrm{W}}$, inclusive cross-section measurements could benefit from an unbinned treatment of the signal process.
@@
The rest of the paper is structured as follows. We discuss the relation to existing works for unbinned SMEFT analyses in Sec. 2. The statistical setup for unbinned hypothesis tests is provided in Sec. 3, and we use it to outline the key concepts of refinable modeling. A comprehensive review of the statistical interpretation of event simulation, suitable for developing ML tools, is given in Sec. 4. This part also defines the terminology for Sec. 5, which explains how to train generic regressors for suitable parametrizations of the various model-parameter dependencies and introduces the BPT algorithm. In Sec. 6, we use the enw tools as building blocks for constructing refinable unbinned models. Those are applied in Sec. 7, where we demonstrate the application of the procedures for SMEFT analyses of top quark pair production in the two-lepton channel. We provide conclusions in Sec. 8.
@@ 2 Relation to other works
The Boosted Information Tree (BIT[10, 9] is a tree-based algorithm for learning SMEFT effects. In this work, we adopt it for the SMEFT signal modeling and extend it toward predicting the full positive quadratic SMEFT polynomial. The BIT algorithm learns the quadratic SMEFT polynomial using the same statistical foundation as the present work. The “Parametric Regression Tree” in Sec. 5.4 has a different goal, but the technical implementation and the statistical interpretation of the boosting are closely related.
@@
The authors of the Madminer framework [13, 14, 15, 16, 17, 18] developed the understanding of event simulation for likelihood-free inference that is also an essential basis for this work. On the technological side, Madminer provides various techniques for general parameter inference and, specifically, also for unbinned SMEFT analyses. Beyond these motivating examples, Madminer arguably established the SBI methods as a subfield in high-energy physics, to which the present work contributes. While Madminer can also model systematic uncertainties, we use a more general and incrementally refinable statistical model for this purpose.
@@ 3 Unbinned likelihood ratio tests
Given a data set $\mathcal{D}$, confidence level (CL) intervals for the POIs ${{\boldsymbol{\theta}}}$ are determined from the profiled likelihood ratio test statistic. In this section, we relate it to quantities that a machine can learn. Splitting the data in a primary set $\mathcal{D}$ and an auxiliary set $\mathcal{A}$, we have
@@
The auxiliary data set has components to constrain systematic uncertainties, such as in the integrated luminosity or the jet energy scale calibration. The maximum-likelihood estimate (MLE) of the nuisance parameters for a given ${{\boldsymbol{\theta}}}$ is $\boldsymbol{\hat{\nu}}_{{\boldsymbol{\theta}}}$, while $(\boldsymbol{\hat{\theta}},\boldsymbol{\hat{\nu}})$ represents the MLE or all model parameters simultaenously.
By design, $q_{{{\boldsymbol{\theta}}}}(\mathcal{D})$ is always non-negative for any ${{\boldsymbol{\theta}}}$, with larger values indicating greater incompatibility between the model defined by ${{\boldsymbol{\theta}}}$ and the data $\mathcal{D}$. Without nuisance parameters, the Neyman-Pearson lemma states that a hypothesis test of fixed size $\alpha$ based on $q_{\boldsymbol{\theta}}(\mathcal{D})$ has maximum power, meaning it is most likely to correctly reject the null hypothesis when the alternative is true [25].
Technically, $\mathcal{A}$ should be an argument of $q_{{{\boldsymbol{\theta}}}}(\mathcal{D})$, but we omit it in the notation, as an analytic approximation of the corresponding likelihood factor, described below, will capture all its effects. We assume that $\mathcal{A}$ and $\mathcal{D}$ do not overlap and that SMEFT effects, parametrized by ${\boldsymbol{\theta}}$, are negligible in $\mathcal{A}$. Under this assumption, the auxiliary data set produces a multiplicative term $L(\mathcal{A}|{\boldsymbol{\nu}})$ in the likelihood,
@@
According to Wilks’ theorem [26], if $\mathcal{D}$ is distributed under $\boldsymbol{\theta}$, then $q_{{{\boldsymbol{\theta}}}}(\mathcal{D})$ is asymptotically distributed as a $\chi^{2}_{N_{{{\boldsymbol{\theta}}}}}$ distribution, where $N_{{{\boldsymbol{\theta}}}}$ is the number of independent degrees of freedom in ${{\boldsymbol{\theta}}}$. This asymptotic distribution is independent of the nuisance parameters, which simplifies the limit-setting procedure and is a primary reason why LHC data analyses commonly use the profiled likelihood ratio test statistic. The confidence intervals at a confidence level of, e.g., $1-\alpha=95\%$, are given by solving
@@
where $F_{\chi^{2}_{N_{{{\boldsymbol{\theta}}}}}}$ is the cumulative distribution function of the $\chi^{2}_{N_{{{\boldsymbol{\theta}}}}}$ distribution.
To obtain confidence intervals with some POIs profiled, these parameters are treated as nuisance parameters in $q_{\boldsymbol{\theta}}$, reducing $N_{{\boldsymbol{\theta}}}$ accordingly. However, large quadratic terms in the SMEFT expansion can invalidate Wilks’ theorem [27]. In such cases, the distribution must be determined by other means, such as toy experiments, to ensure the desired confidence level.
The likelihood functions in Eq. 1 are “extended” likelihoods: a Poisson-distributed counting variable describes the random fluctuation in the total number of observed events. The remaining discriminating information is encoded in the fiducial detector-level probability density function (pdf) $p(\boldsymbol{x}|{{\boldsymbol{\theta}}},{\boldsymbol{\nu}})$, which relates to the fiducial differential cross-section as
@@
We denote the inclusive fiducial cross-section by $\sigma({{\boldsymbol{\theta}}},{\boldsymbol{\nu}})$. In general, ${\textrm{d}}\Sigma(\boldsymbol{x}|{{\boldsymbol{\theta}}},{\boldsymbol{\nu}})$ includes contributions from multiple processes. With the integrated luminosity $\mathcal{L}({\boldsymbol{\nu}})$, subject to systematic uncertainties whose effects we parametrize by nuisance parameters, the likelihood to observe a data set $\mathcal{D}$ of size $N$ can, in general, be written in terms of the differential cross-section as
@@
where $\textrm{P}_{\lambda}$ denotes the Poisson distribution with mean $\lambda$. The extended log-likelihood ratio for two sets of model parameters becomes
@@
Each of the $K$ sources of systematic uncertainty is associated with a nuisance parameter $\nu_{k}$, collectively denoted by ${\boldsymbol{\nu}}$. Systematic effects related to detector calibration, the measurement of the integrated luminosity, theoretical calculations, and more are modeled with these nuisance parameters. Measurements can constrain some of these uncertainties through the observed auxiliary data set $\mathcal{A}_{0}$, which is the specific instance of $\mathcal{A}$ found in real data and, therefore, not a random quantity [22]. We set ${\boldsymbol{\nu}}={\mathbf{0}}$ to correspond to the maximum of $L(\mathcal{A}_{0}|{\boldsymbol{\nu}})$, so that by definition,
@@
The central value ${\boldsymbol{\nu}}={\mathbf{0}}$ represents the best available calibrations across all modeling aspects before considering $\mathcal{D}$. The nuisance parameters then parametrize deviations from this hypothesis. Combined with ${{\boldsymbol{\theta}}}={\mathbf{0}}$, this choice defines an SM reference hypothesis with likelihood
@@
which describes the SM without any SMEFT effects and includes the best available calibrations. We can parametrize the systematic effects to make the $\nu_{k}$ as uncorrelated as possible and scale them so that the auxiliary log-likelihood ratio becomes a simple analytic expression in terms of ${\boldsymbol{\nu}}$. In the Gaussian approximation, this is achieved by diagonalizing the Hessian of the auxiliary likelihood function at ${\boldsymbol{\nu}}={\mathbf{0}}$, resulting in a penalty of the form
@@
though generalizations to other probability distributions are possible.
If a specific nuisance parameter is only constrained by the primary data set and not by $\mathcal{A}_{0}$, it is excluded from the penalty term in Eq. 9 and referred to as “floating”. While some uncertainties, such as those related to PDFs, are clearly interpretable in terms of SM parameters, this is not always the case. For example, uncertainties from renormalization or factorization scales address limitations in perturbative accuracy but do not guarantee statistical coverage when these scales vary in the simulation. With this caveat in mind, we treat all systematic uncertainties heuristically in the same way.
We normalize the likelihoods in Eq. 1 by dividing both the numerator and denominator by the SM reference likelihood in Eq. 8. This yields
$q_{{{\boldsymbol{\theta}}}}(\mathcal{D})=\min_{{\boldsymbol{\nu}}}u(\mathcal{D},\mathcal{A}|{\boldsymbol{\nu}},{{\boldsymbol{\theta}}})-\min_{{\boldsymbol{\nu}},{{\boldsymbol{\theta}}}}u(\mathcal{D},\mathcal{A}|{\boldsymbol{\nu}},{{\boldsymbol{\theta}}}),$ (10)
@@
The main drawback of the unbinned likelihood ratio test statistic in Eq. 11 is the need to evaluate the DCR, inclusive cross-section, and integrated luminosity as functions of the model parameters. While log-normal (multiplicative) nuisances effectively model the integrated luminosity dependence around the central value [28, 29], we also require estimates for the DCR and inclusive cross-section. Event generators cannot provide these estimates parametrically in terms of model parameters, as they operate in “forward” mode through sequential stochastic processes. However, for inference, the minimization in Eq. 1 requires to evaluate the DCR for externally provided simulated or real events.
Figure 2: Overview of refinable modeling in SBI, proceeding in the backward mode, from left to right. For each process, we machine-learn the dependence on the SMEFT POIs and on the nuisance parameters and combine these elements into a refinable surrogate model for the global DCR in the measurement region, suitable for inference.